# ReMoS: 3D Motion-Conditioned Reaction Synthesis for Two-Person Interactions

### v2

@@ Appendix C Additional Results
We provide additional results on the InterHuman [42] dataset, and show how ReMoS can be used as a motion editing tool for character control applications.
C.1 Quantitative Evaluation on the InterHuman Dataset [42]
We report additional evaluation of ReMoS compared to its diffusion-based baselines on the InterHuman [42] dataset in Table 5. We report performance on the standard evaluation metrics, including MPJPE, MPJVE, FID, Diversity and Multi-modality. For Multi-modality, we generate each sequence $5$ times and report numbers with a $95\%$ confidence interval. InterHuman dataset does not provide hand motions, so we only evaluate the reactors’ body motions. ReMoS achieves state-of-the-art performance in the aforementioned metrics in the InterHuman dataset, highlighting the utility of our method for diverse forms of two-person interactions.
C.2 Motion Editing Applications of ReMoS
We describe how to use ReMoS as an interactive motion editing tool, providing control to animators for tasks such as pose completion and motion in-betweening These are crucial applications that are possible due to the strong generative abilities of DDPMs. We provide visual results of these applications in our supplementary video.
@@ C.2 Motion Editing Applications of ReMoS
$X_{B}^{(0)}=\mathit{f}_{{\theta}_{B}}\left(X_{B}^{(t)},t,Y_{B},\mathbbm{1}_{\left\{J_{i},J_{k}\right\}}\right),$ (16)
where $\mathbbm{1}_{\left\{J_{i},J_{k}\right\}}$ is a mask we use at each denoising step on all frames to ensure that the joints $J_{i}$ and $J_{k}$ are not denoised. Instead, we populate $J_{i}$ and $J_{k}$ with the identical noise vectors as the ones used during forward diffusion, while introducing random noise to the rest of the joints throughout the sequence. Thus, animators can incorporate flexible spatial control over chosen joints while ReMoS synthesizes the remaining joints of the reactor to faithfully capture the interaction. In Fig. 5 (bottom left) of the main paper, we show the results of a pose-completion application where we manually control the right-hand wrist joint of the reactor and let ReMoS synthesize the remaining body joints conditioned on the actor.
@@
Likewise, we can use the existing framework to perform motion in-filling for the reactive sequence by providing certain keyframes of the reactive motion, while ReMoS synthesizes the intermediate frames with a motion in-betweening routine. For synthesizing the reactor’s motion between two given keyframes $N_{a}$ and $N_{b}$ through reverse diffusion, we set
$X_{B}^{(0)}=\mathit{f}_{{\theta}_{B}}\left(X_{B}^{(t)},t,Y_{B},\mathbbm{1}_{\left\{N_{a},N_{b}\right\}}\right),$ (17)
where $\mathbbm{1}_{\left\{N_{a},N_{b}\right\}}$ is a mask we use at each denoising step to ensure that all joints at frames $N_{a}$ and $N_{b}$ are not denoised. Thus, ReMoS can fill in the motions between the two seed frames as shown in Fig. 5 (bottom right) of the main paper.

### v3

@@ 0.D.1 Quantitative Evaluation on the InterHuman Dataset [42]
Table 0.D.2: Trainable parameter counts.
Method Params
(full model)
InterFormer $8.2M$
MixNMatch $\mathbf{6.5M}$
ComMDM $22.2M$
RAIG $81.2M$
InterGen $170M$
ReMoS (ours) $17.4M$
Table 0.D.3: Quantitative evaluation on body joints. We compare the body synthesis module of ReMoS with state-of-the-art motion synthesis methods on body joints only. Bold indicates the best.
Methods Lindy Hop (body only) Ninjutsu (body only)
MPJPE $\downarrow$ MPJVE $\downarrow$ FID $\downarrow$ Div $\rightarrow$ MPJPE $\downarrow$ MPJVE $\downarrow$ FID $\downarrow$ Div $\rightarrow$
GT - - - $7.62$ $11.5$
MixNMatch $69.8$ $10.5$ $0.74$ $2.52$ $260.1$ $5.14$ $0.72$ $4.94$
InterFormer $63.2$ $8.91$ $0.52$ $4.64$ $262.5$ $3.53$ $0.51$ $6.27$
ComMDM $50.2$ $4.42$ $0.23$ $7.51$ $192.4$ $3.45$ $0.25$ $9.83$
RAIG $68.5$ $4.01$ $0.26$ $9.02$ $188.3$ $4.25$ $0.19$ $10.14$
InterGen $55.1$ $2.87$ $0.22$ $7.49$ $165.5$ $3.82$ $0.23$ $9.87$
ReMoS (ours) $\mathbf{40.2}$ $\mathbf{2.21}$ $\mathbf{0.12}$ $\mathbf{7.52}$ $\mathbf{137.2}$ $\mathbf{3.19}$ $\mathbf{0.16}$ $\mathbf{10.26}$
0.D.2 Trainable Parameters
We report the total number of trainable parameters of ReMoS as compared to the baseline methods. Table 0.D.2 shows that ReMoS has lesser trainable parameters than the existing diffusion-based two-person synthesis models [54, 65, 42].
0.D.3 Comparison with baselines without hand motions.
We compare the body synthesis module of ReMoS with baselines trained only on the body joints (Table 0.D.3). We report state-of-the-art performance for ReMoS even when finger joints are not included.
0.D.4 Motion Editing Applications of ReMoS
We describe how to use ReMoS as an interactive motion editing tool, providing control to animators for tasks such as pose completion and motion in-betweening. These are crucial applications that are possible due to the strong generative abilities of DDPMs. We provide visual results of these applications in our supplementary video.
@@ 0.D.4 Motion Editing Applications of ReMoS
$X_{B}^{\left(0\right)}=\mathit{f}_{{\theta}_{B}}\left(X_{B}^{\left(t\right)},t,Y_{B},\mathbbm{1}_{\left\{J_{i},J_{k}\right\}}\right),$ (0.D.1)
where $\mathbbm{1}_{\left\{J_{i},J_{k}\right\}}$ is a mask we use at each denoising step on all frames to ensure that the joints $J_{i}$ and $J_{k}$ are not denoised. Instead, we populate $J_{i}$ and $J_{k}$ with the identical noise vectors as the ones used during forward diffusion, while introducing random noise to the rest of the joints throughout the sequence. Thus, animators can incorporate flexible spatial control over chosen joints while ReMoS synthesizes the remaining joints of the reactor to faithfully capture the interaction. In Fig. 5b in the main paper, we show the results of a pose-completion application where we manually control the right-hand wrist joint of the reactor and let ReMoS synthesize the remaining body joints conditioned on the actor.
@@
Likewise, we can use the existing framework to perform motion in-betweening for the reactive sequence. We achieve this by providing some keyframes of the reactive motion and letting ReMoS synthesize the intermediate frames using a motion in-betweening routine. To synthesize the reactor’s motion between two given keyframes $N_{a}$ and $N_{b}$ through reverse diffusion, we set
$X_{B}^{\left(0\right)}=\mathit{f}_{{\theta}_{B}}\left(X_{B}^{\left(t\right)},t,Y_{B},\mathbbm{1}_{\left\{N_{a},N_{b}\right\}}\right),$ (0.D.2)
where $\mathbbm{1}_{\left\{N_{a},N_{b}\right\}}$ is a mask we use at each denoising step to ensure that all joints at frames $N_{a}$ and $N_{b}$ are not denoised. Thus, ReMoS can fill in the motions between the two seed frames as shown in Fig. 5c in the main paper.